(1) Suppose you have a rectangle with length 360 units and width 84 units. Each turn, you cut off the greatest possible square from the rectangle. You do so until you have only squares. How many squares will you get?

Answer these two other questions Please!

(2) n is a natural number. Find the greatest possible value of the GCD of
(2n+25) and (n+15)

(3) n is a natural number. Find the greatest possible value of the GCD of
2(n+8) and (n+13)
NEED ASAP!

Answer by ikleyn(52781)

.


I will help you to solve the second problem.

    n is a natural number.  Find the common Greatest Common Divisor (GCD) of numbers (2n+25) and (n+15).


                        Solution


Let integer number d is the common divisor of (2n+25) and (n+5).


Then it divides the difference 


    (2n+25) - 2*(n+5) = (2n+5) - (2n+10) = 2n + 5 - 2n - 1- = -5.


Thus, any common divisor of the numbers (2n+5) and (n+5) divides the number -5.


It means that the Common Greatest Divisor (GCD) of these numbers is EITHER 5 OR 1.


It can not be any other number.




Example 1:   Let n = 10.

             Then the numbers (2n+5) = 25  and  (n+5) = 15.

             Their Greatest Common Divisor is 5.



Example 2:   Let n = 7.

             Then the numbers (2n+5) = 19  and  (n+5) = 12.

             Their Greatest Common Divisor is 1.



Thus I explained you the solution and illustrated it by examples.


ANSWER.    GCD of these numbers is   EITHER  5  OR  1.


The solution is completed.


After learning it from me,  now solve the  problem #3  ON  YOUR  OWN.


/\/\/\/\/\/\/\/


NEVER  submit more than one problem per post.

It is the  RULE,  the  REQUIREMENT  and the  POLICE  of this forum.

It is the  NECESSARY  condition to communicate.

Do not violate it and do not work against yourself.